OpenStudy (anonymous):

WILL FAN + MEDAL :) If the following is a polynomial function, then state its degree and leading coefficient. If it is not, then state this fact. f(x) = -16x5 - 7x4 - 6 A Degree: -16; leading coefficient: 5 B Degree: 5; leading coefficient: -16 C Not a polynomial function. D Degree: 9; leading coefficient: -16

OpenStudy (anonymous):

Is it B?

OpenStudy (dayakar):

correct,

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@dayakar @terron_jones2003 just making sure thanks! Can you check if i did one more thing right :)

OpenStudy (anonymous):

Degree: 5; leading coefficient: 10

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Write the quadratic function in vertex form. y = x2 + 4x + 7 is the answer y = (x + 2)^2 + 3

OpenStudy (anonymous):

y =(x + 2)^2 + 3. Your right! Good Job!

OpenStudy (dayakar):

exactly,well done

OpenStudy (anonymous):

Thanks! :)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

no problem!

OpenStudy (dayakar):

welcome

OpenStudy (anonymous):

can i get a medal for helping